# Large-sample analysis of cost functionals for inference under the coalescent

### v1

@@ 7.1.1 Technical framework and additional notation
For the limiting process $\mathbf{Z}$, we thus consider the state space $E=\mathbb{R}_{+}\times E_{1}\times\mathbb{N}^{d^{2}}$, where $E_{1}=(0,\infty]^{d}$. We equip $E$ with the product metric $\psi=\left\lVert\cdot\right\rVert_{2}\oplus\psi_{1}\oplus\left\lVert\cdot\right\rVert_{2}$, where $\psi_{1}(\mathbf{y}_{1},\mathbf{y}_{2})=\left\lVert 1/\mathbf{y}_{1}-1/\mathbf{y}_{2}\right\rVert_{2}$, with component-wise inversion and with the inverse of $\infty$ being 00. Note that, in $E_{1}$, the roles of 00 and $\infty$ are reversed component-wise, the metric $\psi_{1}$ is equivalent to the Euclidean metric away from the boundary $\Omega_{0}$ and from infinity, and compact sets are bounded away from $\Omega_{0}$.
Let $C_{c}^{\infty}(E)$ and $\hat{C}(E)$ be the spaces of real-valued continuous functions on $(E,\psi)$ that are, respectively, smooth with compact support or vanishing at infinity. In $(E,\psi)$, functions with compact support are equal to zero near $\Omega_{0}$ in the $E_{1}$-component and near the classical infinity, in the other components. Similarly, functions vanishing at infinity, vanish towards $\Omega_{0}$ in the $E_{1}$-component and towards infinity, in the classical sense, in the other components. For further explanations and properties of state spaces and related functions we refer to (Favero and Hult, 2024, Appendix A.2).
Furthermore, let $E^{(n)}=\mathbb{R}_{+}\times\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}\times\mathbb{N}^{d^{2}}$ be the state space of $\mathbf{Z}^{(n)}$, and let $\eta_{n}$ map any function on $E$ into its restriction on $E^{(n)}$, with value zero on +×$\mathbb{R}_{+}\times\Omega_{0}\times\mathbb{N}^{d^{2}}$.
@@ 7.1.2 Convergence of generators (PIM)
$K=\{(c,\mathbf{y},\mathbf{m})\in E:y_{i}\geq\delta,c\leq M,m_{ij}\leq M,\forall i,j=1,\dots,d\}.$
@@
$\lim_{n\to\infty}\sup_{\mathbf{y}\in E_{1}^{(n)}\cap K_{1}}\left|n(c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-1)-a_{j}(\mathbf{y})\right|=0,\qquad\lim_{n\to\infty}\sup_{\mathbf{y}\in E_{1}^{(n)}\cap K_{1}}\left|c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y})-b_{ij}(\mathbf{y})\right|=0,$ (7.2)
for $i,j=1,\dots,d$. Furthermore, in (Favero and Hult, 2024, Proof of Theorem 2.1) it is shown, in the PIM case, that
$\lim_{n\to\infty}\sup_{\mathbf{y}\in E_{1}^{(n)}\cap K_{1}}\left|\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y})-\frac{y_{j}}{\|\mathbf{y}\|_{1}}\right|=0,\qquad\lim_{n\to\infty}\sup_{\mathbf{y}\in E_{1}^{(n)}\cap K_{1}}\left|n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})-\lambda_{ij}(\mathbf{y})\right|=0,$ (7.3)
for $i,j=1,\dots,d$.
To prove (7.1), we first take $(c,\mathbf{y},\mathbf{m})\in E^{(n)}\cap K^{\complement}$. Then, $f=Af=0$ in a neighbourhood of $(c,\mathbf{y},\mathbf{m})$. If also $\left(c\,c^{(n)}(\mathbf{e}_{j}|\mathbf{y}),\mathbf{y}^{(n)}-\frac{1}{n}\mathbf{e}_{j},\mathbf{m}\right)$ and $\left(c\,c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j}+\frac{1}{n}\mathbf{e}_{i},\mathbf{m}+\mathbf{e}_{ij}\right)$ belong to $E^{(n)}\cap K^{\complement}$, for all $i,j=1,\dots,d,n\in\mathbb{N}$, then $A^{(n)}\eta_{n}f(c,\mathbf{y},\mathbf{m})=\eta_{n}Af(c,\mathbf{y},\mathbf{m})=0.$ Otherwise, it must be that $m_{ij}<M,i,j=1,\dots,d$, and one of the following two cases occurs:
@@
    For a unique $i_{0}$ and some $n$, $\delta-1/n\leq y_{i_{0}}<\delta$, while $y_{j}\geq\delta$ for all $j\neq i_{0}$ and $c\leq M$, $cc^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})\leq M$, $cc^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y})\leq M$, $i,j=1,\dots,d$;
@@
    $y_{j}\geq\delta$ for all $j=1,\dots,d$, $c>M$, and, for some $j$ and/or $i$, $cc^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})\leq M$, and/or $cc^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y})\leq M$.
In both cases, $A^{(n)}\eta_{n}f(c,\mathbf{y},\mathbf{m})$ is different from zero, but converges uniformly to 00 because $\mathbf{y}\in K_{1}$, $b_{ij}\geq 1,i,j=1,\dots,d$, and because of (7.2), (7.3), and the properties of $f$.
Now, we take $(c,\mathbf{y},\mathbf{m})\in E^{(n)}\cap K$ and find a bound for $\left|A^{(n)}\eta_{n}f(c,\mathbf{y},\mathbf{m})-\eta_{n}Af(c,\mathbf{y},\mathbf{m})\right|$. First, note that, for $j=1,\dots,d,$ there exist $\bar{c}_{j}$, with $|\bar{c}_{j}-c|\leq|cc^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-c|$, and $s_{j}$, with $|s_{j}|\leq 1/n$, such that
@@
$\qquad=\partial_{c}f(\bar{c}_{j},\mathbf{y}-s_{j}\mathbf{e}_{j},\mathbf{m})c(c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-1)-\frac{1}{n}\partial_{y_{j}}f(\bar{c}_{j},\mathbf{y}-s_{j}\mathbf{e}_{j},\mathbf{m}).$
@@
$\qquad+\sum_{j=1}^{d}\left|\partial_{y_{j}}f(\bar{c}_{j},\mathbf{y}-s_{j}\mathbf{e}_{j},\mathbf{m})\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y})-\partial_{y_{j}}f(c,\mathbf{y},\mathbf{m})\frac{y_{j}}{\|\mathbf{y}\|_{1}}\right|$ (7.5)
$\qquad+\sum_{i,j=1}^{d}\Bigg{|}f\left(c\,c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j}+\frac{1}{n}\mathbf{e}_{i},\mathbf{m}+\mathbf{e}_{ij}\right)n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})-$
$\qquad\quad\quad\quad-f(cb_{ij}(\mathbf{y}),\mathbf{y},\mathbf{m}+\mathbf{e}_{ij})\lambda_{ij}(\mathbf{y})\Bigg{|}$ (7.6)
$\qquad+|f(c,\mathbf{y},\mathbf{m})|\sum_{i,j=1}^{d}|n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})-\lambda_{ij}(\mathbf{y})|.$ (7.7)
@@
$\qquad+M\left\lVert\partial_{c}f\right\rVert_{\infty}\left|n(c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-1)\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y})-a_{j}(\mathbf{y})\frac{y_{j}}{\|\mathbf{y}\|_{1}}\right|,$
@@
$\qquad+\left\lVert f\right\rVert_{\infty}\left|n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})-\lambda_{ij}(\mathbf{y})\right|,$
@@ 7.1.3 Weak convergence (general mutation)
To prove the result in the general mutation case, we can use the change-of-measure argument developed in (Favero and Hult, 2024, Sect. 3). This consists of changing the measures so that, under the new measures, the originally parent-dependent mutations become parent-independent. Crucially, the Radon-Nykodym derivatives (likelihood ratios) of the changes of measure depend on the block-counting and mutation-counting components, $\mathbf{Y}^{(n)},\mathbf{Y},\mathbf{M}^{(n)},\mathbf{M}$, not on the cost-counting components, $C^{(n)},C$, and thus are exactly the same as in (Favero and Hult, 2024), where the cost-components are not considered. Then, the PIM results can be applied to complete the proof in the general case, see (Favero and Hult, 2024, Sect. 5.4) for details.
@@ 7.5.1 Griffiths–Tavaré
[…]
@@ 7.5.2 Stephens–Donnelly
[…]
@@ 8 Discussion
In genetics, the coalescent is a robust model for a wide range of settings and organisms, but relies on a small variance of family sizes relative to population size. If family sizes are heavily skewed, evolution can be more accurately described by multiple merger coalescents, in which more than two lineages can coalesce simultaneously (Donnelly and Kurtz, 1999; Pitman, 1999; Sagitov, 1999), and more than one simultaneous coalescence (Möhle and Sagitov, 2001; Schweinsberg, 2000) can take place. Importance sampling methods for these types of models are available but are even less scalable as those for the standard coalescent (Birkner et al., 2011; Koskela et al., 2015). A similar scaling limit for multiple merger coalescents would be of mathematical interest, and could inform importance sampling methods for them as well. If such a scaling limit exists, we expect it would incorporate macroscopic jumps in towards the origin driven by multiple mergers.
Finally, modern data sets rarely consist of a single locus. Hence it would be of interest to obtain a similar description of weighted ancestral recombination graphs, which are the multi-locus analogue of the coalescent. Evolution at two unlinked loci would correspond to two independent copies of our limiting process. A scaling limit for two linked loci should be informative of how linkage creates correlation between the two copies of the limit process. Such a result would be of mathematical interest, and could also inform Monte Carlo methods (Fearnhead and Donnelly, 2001) and more heuristic methods (Li and Stephens, 2003) for genomic inference.

### v2

@@ 7.1.1 Technical framework and additional notation
For the limiting process $\mathbf{Z}$, we thus consider the state space $E=\mathbb{R}_{+}\times E_{1}\times\mathbb{N}^{d^{2}}$, where $E_{1}=(0,\infty]^{d}$. We equip $E$ with the product metric $\psi=\left\lVert\cdot\right\rVert_{2}\oplus\psi_{1}\oplus\left\lVert\cdot\right\rVert_{2}$, where $\psi_{1}(\mathbf{y}_{1},\mathbf{y}_{2})=\left\lVert 1/\mathbf{y}_{1}-1/\mathbf{y}_{2}\right\rVert_{2}$, with component-wise inversion and with the inverse of $\infty$ being 0. Note that, in $E_{1}$, the roles of 0 and $\infty$ are reversed component-wise, the metric $\psi_{1}$ is equivalent to the Euclidean metric away from the boundary $\Omega_{0}$ and from infinity, and compact sets are bounded away from $\Omega_{0}$.
Let $C_{c}^{\infty}(E)$ and $\hat{C}(E)$ be the spaces of real-valued continuous functions on $(E,\psi)$ that are, respectively, smooth with compact support or vanishing at infinity. In $(E,\psi)$, functions with compact support are equal to zero near $\Omega_{0}$ in the $E_{1}$-component and near the classical infinity, in the other components. Similarly, functions vanishing at infinity, vanish towards $\Omega_{0}$ in the $E_{1}$-component and towards infinity, in the classical sense, in the other components. For further explanations and properties of this state space and related functions we refer to (Favero and Hult, 2024, Appendix A.2).
Furthermore, let $E^{(n)}=\mathbb{R}_{+}\times\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}\times\mathbb{N}^{d^{2}}$ be the state space of $\mathbf{Z}^{(n)}$, and let $\eta_{n}$ map any function on $E$ into its restriction on $E^{(n)}$, with value zero on $\mathbb{R}_{+}\times\Omega_{0}\times\mathbb{N}^{d^{2}}$.
@@ 7.1.2 Convergence of generators (PIM)
$K=\{(c,\mathbf{y},\mathbf{m})\in E:y_{i}\geq\delta,c\leq M,m_{ij}\leq M,\forall i,j\in\{1,\dots,d\}\}.$
@@
$\lim_{n\to\infty}\sup_{\mathbf{y}\in E_{1}^{(n)}\cap K_{1}}\left|n(c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-1)-a_{j}(\mathbf{y})\right|=0,\hskip 18.49988pt\lim_{n\to\infty}\sup_{\mathbf{y}\in E_{1}^{(n)}\cap K_{1}}\left|c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y})-b_{ij}(\mathbf{y})\right|=0,$ (7.2)
for $i,j\in\{1,\dots,d\}$. Furthermore, in (Favero and Hult, 2024, Proof of Theorem 2.1) it is shown, in the PIM case, that
$\lim_{n\to\infty}\sup_{\mathbf{y}\in E_{1}^{(n)}\cap K_{1}}\left|\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y})-\frac{y_{j}}{\|\mathbf{y}\|_{1}}\right|=0,\hskip 18.49988pt\lim_{n\to\infty}\sup_{\mathbf{y}\in E_{1}^{(n)}\cap K_{1}}\left|n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})-\lambda_{ij}(\mathbf{y})\right|=0,$ (7.3)
for $i,j\in\{1,\dots,d\}$.
To prove (7.1), we first take $(c,\mathbf{y},\mathbf{m})\in E^{(n)}\cap K^{\complement}$. Then, $f=Af=0$ in a neighbourhood of $(c,\mathbf{y},\mathbf{m})$. If also $\left(c\,c^{(n)}(\mathbf{e}_{j}|\mathbf{y}),\mathbf{y}^{(n)}-\frac{1}{n}\mathbf{e}_{j},\mathbf{m}\right)$ and $\left(c\,c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j}+\frac{1}{n}\mathbf{e}_{i},\mathbf{m}+\mathbf{e}_{ij}\right)$ belong to $E^{(n)}\cap K^{\complement}$, for all $i,j\in\{1,\dots,d\}$, $n\in\mathbb{N}$, then $A^{(n)}\eta_{n}f(c,\mathbf{y},\mathbf{m})=\eta_{n}Af(c,\mathbf{y},\mathbf{m})=0.$ Otherwise, it must be that $m_{ij}<M,i,j\in\{1,\dots,d\}$, and one of the following two cases occurs:
@@
    For a unique $i_{0}$ and some $n$, $\delta-1/n\leq y_{i_{0}}<\delta$, while $y_{j}\geq\delta$ for all $j\neq i_{0}$ and $\leq M, cc$$^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})\leq M, cc$$c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y})\leq M$, $i,j\in\{1,\dots,d\}$;
@@
    $y_{j}\geq\delta$ for all $j\in\{1,\dots,d\}$, $c>M$, and, for some $j$ and/or $i$, $cc^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})\leq M$, and/or $cc^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y})\leq M$.
In both cases, $A^{(n)}\eta_{n}f(c,\mathbf{y},\mathbf{m})$ is different from zero, but converges uniformly to 0 because $\mathbf{y}\in K_{1}$, $b_{ij}\geq 1,\;i,j\in\{1,\dots,d\}$, and because of (7.2), (7.3), and the properties of $f$.
Now, we take $(c,\mathbf{y},\mathbf{m})\in E^{(n)}\cap K$ and find a bound for $\left|A^{(n)}\eta_{n}f(c,\mathbf{y},\mathbf{m})-\eta_{n}Af(c,\mathbf{y},\mathbf{m})\right|$. First, note that, for $j\in\{1,\dots,d\},$ there exist $\bar{c}_{j}$, with $|\bar{c}_{j}-c|\leq|cc^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-c|$, and $s_{j}$, with $|s_{j}|\leq 1/n$, such that
@@
$\hskip 18.49988pt=\partial_{c}f(\bar{c}_{j},\mathbf{y}-s_{j}\mathbf{e}_{j},\mathbf{m})c(c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-1)-\frac{1}{n}\partial_{y_{j}}f(\bar{c}_{j},\mathbf{y}-s_{j}\mathbf{e}_{j},\mathbf{m}).$
@@
$\hskip 18.49988pt+\sum_{j=1}^{d}\left|\partial_{y_{j}}f(\bar{c}_{j},\mathbf{y}-s_{j}\mathbf{e}_{j},\mathbf{m})\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y})-\partial_{y_{j}}f(c,\mathbf{y},\mathbf{m})\frac{y_{j}}{\|\mathbf{y}\|_{1}}\right|$ (7.5)
$\hskip 18.49988pt+\sum_{i,j=1}^{d}\Bigg|f\left(c\,c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j}+\frac{1}{n}\mathbf{e}_{i},\mathbf{m}+\mathbf{e}_{ij}\right)n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})$-
$\hskip 18.49988pt\qquad\qquad\qquad-f(cb_{ij}(\mathbf{y}),\mathbf{y},\mathbf{m}+\mathbf{e}_{ij})\lambda_{ij}(\mathbf{y})\Bigg|$ (7.6)
$\hskip 18.49988pt+|f(c,\mathbf{y},\mathbf{m})|\sum_{i,j=1}^{d}|n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})-\lambda_{ij}(\mathbf{y})|.$ (7.7)
@@
$\hskip 18.49988pt+M\left\lVert\partial_{c}f\right\rVert_{\infty}\left|n(c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-1)\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y})-a_{j}(\mathbf{y})\frac{y_{j}}{\|\mathbf{y}\|_{1}}\right|,$
@@
$\hskip 18.49988pt+\left\lVert f\right\rVert_{\infty}\left|n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})-\lambda_{ij}(\mathbf{y})\right|,$
@@ 7.1.3 Weak convergence (general mutation)
To prove the result in the general mutation case, we can use the change-of-measure argument developed in (Favero and Hult, 2024, Sect. 3). This consists of changing the measures so that, under the new measures, the originally parent-dependent mutations become parent-independent. Crucially, the Radon–Nikodym derivatives (likelihood ratios) of the changes of measure depend on the block-counting and mutation-counting components, $\mathbf{Y}^{(n)},\mathbf{Y},\mathbf{M}^{(n)},\mathbf{M}$, not on the cost-counting components, $C^{(n)},C$, and thus are exactly the same as in (Favero and Hult, 2024), where the cost-components are not considered. Then, the PIM results can be applied to complete the proof in the general case, see (Favero and Hult, 2024, Sect. 5.4) for details.
@@ 7.5.1 Griffiths–Tavaré
$C^{(n)}_{GT}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}C_{GT}(t)=$ $\exp\left\{\sum_{i=1}^{d}y_{0,i}\int_{0}^{t}\frac{1-d}{\|\mathbf{y}_{0}(1-u)\|_{1}}du\right\}$
$=$ $\exp\left\{\int_{0}^{t}\frac{1-d}{1-u}du\right\}$
@@ 7.5.2 Stephens–Donnelly
$C^{(n)}_{SD}(\lfloor{tn}\rfloor{})\xrightarrow[n\to\infty]{\mathcal{D}}C_{SD}(t)=\exp\left\{\sum_{i=1}^{d}y_{0,i}\int_{0}^{t}\hat{a}_{i}\left(\mathbf{y}_{0}(1-u)\right)du\right\}=\exp\left\{\int_{0}^{t}\frac{1-d}{1-u}du\right\}=(1-t)^{d-1},$
@@ 8 Discussion
In genetics, the coalescent is a robust model for a wide range of settings and organisms, but relies on a small variance of family sizes relative to population size. If family sizes are heavily skewed, evolution can be more accurately described by multiple merger coalescents, in which more than two lineages can coalesce simultaneously (Donnelly and Kurtz, 1999; Pitman, 1999; Sagitov, 1999), and more than one simultaneous coalescence can take place (Möhle and Sagitov, 2001; Schweinsberg, 2000). Importance sampling methods for these types of models are available but are even less scalable as those for the standard coalescent (Birkner and Blath, 2008; Birkner et al., 2011; Koskela et al., 2015). A similar scaling limit for multiple merger coalescents would be of mathematical interest, and could inform importance sampling methods for them as well. If such a scaling limit exists, we expect it would incorporate macroscopic jumps towards the origin driven by multiple mergers.
Finally, modern data sets rarely consist of a single locus. Hence it would be of interest to obtain a similar description of weighted ancestral recombination graphs, which are the multi-locus analogue of the coalescent. Evolution at two unlinked loci would correspond to two independent copies of our limiting process. A scaling limit for two linked loci should be informative of how linkage creates correlation between the two copies of the limit process. Such a result would be of mathematical interest, and could also inform importance sampling methods (Fearnhead and Donnelly, 2001) and more heuristic methods (Li and Stephens, 2003) for genomic inference.